One-sixth of a number is 5 more than the same number. What is the number?
Answer by checkley77(12569)

x/6=x+5
x=6(x+5)
x=6x+30
5x=-30
x=-30/5
x=-6 ans.
Proof:
-6/6=-6+5
-1=-1